1A) Write a balanced equation for the following reaction in basic solution.


Fe(OH)3 + ClO → FeO42−+Cl

1B) Use VSEPR theory to predict a molecular shape for XeF2 and hence determine the total number of vibrational modes expected to be observed in its IR and Raman spectra. Would any of these absorptions occur at the same frequency in both Raman and IR spectra?
Answer #1

Reduction 1/A) 2- Te GHz + do Feo-tu- oridation Staty of non (+3) X+62)=-1 X+4(2)=-2 -1 Fe, el x = 1 x = 6 1 Orédation state